find the equation tot he straight line passing through the point (3,2)and the point of intersection of the lines 2x+3y=1 and 3x-4y=6

Grade:11

1 Answers

Utkarsh Verma
10 years ago

First, find the common point between the two intersecting lines,

2x+3y=1  and  3x-4y=6 .

Solve them as a set of equations and you will get  x = 22/7 and y = -9/17

Then find the slope using, m = (y2-y1) / (x2-x1)

You''ll find, m = 43 / 29

Use y - y0 = m (x - x0) to get the equation.

It will look like: y - 2 = (43 / 29) (x - 3)

 

 


 

The final equation will be: 43x - 29y - 71 = 0